A 700 W lamp was connected to a supply that a 5A current was flowing through the lamp. What is that resistance of the lamp
1
Expert's answer
2021-09-21T11:05:07-0400

power, "P=I^2R"

"R=\frac{P}{I^2}=\frac{700}{5^2}=" "28\Omega"